# Refining Neural Activation Patterns for Layer-Level Concept Discovery in Neural Network-Based Receivers

### v1

@@ 2 Methodology
where $K()$ is the kernel function (e.g., Gaussian), $H$ is a smoothing parameter called the bandwidth, and $Q$ is the number of activation values on channel $u$ in layer $l$. We use Scott’s Rule (Scott (1992)) for selecting the bandwidth, providing a standard bias-variance tradeoff:
$\phantom{.}H=1/\sqrt[5]{Q}.$ (3)
where $U$ is the number of channels in layer $l$, using HDBSCAN* with the Jensen-Shannon distance metric (Lin (1991); Endres and Schindelin (2003)), which is symmetric, bounded, and well-suited for comparing probability distributions such as those estimated by KDE in the previous step. The Jensen-Shannon distance $D^{l}_{\text{JS}}$ between activations for inputs $x_{i}$ and $x_{j}$ for all channels in layer $l$ is
$D^{l}_{\text{JS}}(x_{i},x_{j})=\frac{1}{\sqrt{2}}\sqrt{\sum^{U\cdot R}_{k=1}{\left(\mathbf{\hat{g}}^{l}_{i,k}\log\frac{2\cdot\mathbf{\hat{g}}^{l}_{i,k}}{\mathbf{\hat{g}}^{l}_{i,k}+\mathbf{\hat{g}}^{l}_{j,k}}+\mathbf{\hat{g}}^{l}_{j,k}\log\frac{2\cdot\mathbf{\hat{g}}^{l}_{j,k}}{\mathbf{\hat{g}}^{l}_{i,k}+\mathbf{\hat{g}}^{l}_{j,k}}\right)}}$ (7)
and we use this as a distance metric in clustering.
Proposed Post-Clustering Analysis: We evaluate clustering quality using two internal metrics and one application-specific metric. As internal metrics, we use the persistence score (Campello et al. (2013)), which reflects the stability or robustness of a cluster (with higher values indicating more well-defined clusters), and the Density-Based Clustering Validation (DBCV) index (Moulavi et al. (2014)), which is designed to assess density-based clustering algorithms like HDBSCAN by balancing cluster compactness and separation. As an application-specific metric, we use Out-of-Distribution (OOD) detection (see, e.g. Zhu et al. (2025)), assuming that OOD samples fall outside the learned cluster structure. This allows clustering to act as a proxy for distributional mismatch. In this setting clustering serves not just as an explanation method but as a model debugging tool—similar to how Adebayo et al. (2020) frame OOD behavior as test-time contamination. To interpret the resulting clusters, we use the channel parameters of the inputs (e.g., SNR) to summarize and explain structure.
3 Experiments and Results on Visual Object Recognition Model Activations
We applied the NAP methodology using the average amount of activations as the baseline, as suggested by Bäuerle et al. (2022) to be the most suitable choice for extracting a larger number of NAPs. We then compared this baseline with our proposed methodology—outlined in Figure 2 and detailed in Section 2—and additionally studied the effect of HDBSCAN parameters.
3.1 Experimental Setup
The deep neural network-based visual object classification model was based on the ResNet-50 architecture He et al. (2016), using a pretrained weights from TorchVision (Marcel and Rodriguez (2010)). Further implementation and dataset details are provided in Appendix C.
We used our own PyTorch-based implementation of both the NAP baseline and our proposed methodology; implementation details are provided in Appendix B.
3.2 Results
We first evaluated the impact of dataset size using the HDBSCAN parameters recommended by Bäuerle et al. (2022)—specifically, the Leaves of Tree (Leaf) as the cluster selection criteria with a minimum cluster size of 5. As shown in Figure 4, the baseline method resulted in 73–95% of samples remaining unclustered. Our proposed methodology did not mitigate this issue. A high proportion of non-clustered activations is problematic, as it hinders the discovery of meaningful layer-level concepts, reducing interpretability. It also undermines confidence in the clustering structure and impairs downstream tasks such as Out-of-Distribution (OOD) detection, which depend on well-formed, dense regions in activation space.
Notably, Bäuerle et al. (2022) do not report any analysis of unclustered or outlier samples, which we found to be a recurring issue in the visual object recognition setting as well. Further results (Appendix E) show that increasing dataset size leads to more clusters being discovered—consistent with findings in Bäuerle et al. (2022)—but the mean persistence score remains relatively stable.
As our primary focus is on analyzing activations from the radio receiver model, we conclude our analysis of the visual object recognition results here and proceed to the next set of experiments.
4 Experiments and Results on Radio Receiver Model Activations
Following the approach used in Section 3, we applied both the baseline and proposed methodologies—illustrated in Figure 2 and described in detail in Section 2 and in Appendix A—to a generated dataset. Our goal was to analyze and summarize the layer-level concepts learned by the deep neural network-based receiver model in terms of underlying channel parameters.
4.1 Experimental Setup
The deep neural network-based receiver model (DeepRx, as detailed in Appendix D) was based on Honkala et al. (2021), adopting the Performer Model from Table 1 of Tuononen et al. (2025), which presents a more minimal version. The model was trained following Honkala et al. (2021). Training and testing data were generated using link-level simulations with the NVIDIA Sionna library (Hoydis et al. (2023)), as detailed in Table 1 in Appendix D. We applied the same NAP baseline and proposed methodology as described in Section 3.1, with implementation details provided in Appendix B.
4.2 Clustering Sensitivity to Parameters and Preprocessing
We first evaluated the impact of dataset size using HDBSCAN with the Leaf cluster selection criterion and a minimum cluster size of 5, as recommended by Bäuerle et al. (2022). As shown in Figure 6, the baseline method resulted in 59–97% of samples remaining unclustered, with the proportion increasing alongside dataset size. Our proposed methodology did not resolve this issue. As in the previous experiment (Section 3.2), more data led to more clusters being discovered (Appendix F), but here the mean persistence score tended to decrease.
We then explored the effect of HDBSCAN parameters. As shown in Figure 6, using the EoM criterion and smaller minimum cluster sizes (e.g., 3 or 5) consistently reduced the number of unclustered samples. Our proposed method generally offered more stable performance, avoiding sharp changes seen in the baseline (e.g., at cluster size 13 on B1-PRE, or high unclustered rates on OUT-POST). Cluster count trends mirrored those in the ImageNet experiment, except for OUT-POST, where both criteria showed decreasing cluster counts with increasing minimum cluster size.
Applying Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction with standard settings before HDBSCAN (Appendix L) generally increased both the number of clusters and unclustered samples, making interpretation more difficult. Although a few configurations (e.g., proposed method with EoM, minimum cluster size 19, and 15 UMAP dimensions on B1-PRE) reduced unclustered samples to near-zero, these effects were inconsistent across layers and methodologies. We suspect that UMAP may distort the data structure in ways that occasionally aid clustering, but its behavior was too unstable to be useful in practice (Appendix I).
Finally, applying HDBSCAN directly to raw activations—i.e., without aggregation or normalization—left all samples unclustered (see Appendices J and L), likely due to the curse of dimensionality (Bellman (1961)). Adding UMAP for dimensionality reduction in this setup sometimes enabled full clustering—for example, using EoM as the cluster selection criterion with a minimum cluster size of 3 on layer B1-PRE and 15 UMAP dimensions—but results remained erratic and inconsistent across layers and parameters. As such, we do not further consider the "no aggregation and no normalization" methodology or UMAP-based dimensionality reduction in this work.
Figure 5: Percentage of clustered and non-clustered samples as function of sample size and aggregation method with minimum cluster size 5 using link-level simulations with DeepRx model.
Figure 6: Number of non-clustered samples and clusters by aggregation and cluster selection method as function of minimum cluster size using link-level simulations with DeepRx model.
4.3 Layer-Level Concepts in Terms of Channel Parameters
We found no strong clustering structure in the data. When HDBSCAN parameters were tuned to minimize unclustered samples, nearly all data fell into a single large cluster (Figure 6; bottom of Figure 7). Internal clustering quality metrics—DBCV and persistence scores—were consistently near zero (Appendix H), indicating weak and unstable groupings. Most samples formed one incoherent cluster, with remaining clusters small and likely spurious.
In contrast, 2D UMAP embeddings revealed a continuous manifold shaped by SNR (Figure 7, top). Under the baseline, data followed a single SNR-driven path. The proposed method produced two distinct paths that converge at low SNR, forming a unified structure. SNR clearly dominates the activation space, with the channel model introducing variation at high SNR and minor deviations from the main manifold among high delay spread samples. The dominance of SNR—and the limited influence of UE speed and delay spread—aligns with our replication using those parameters (Appendix G) with methodology of Tuononen et al. (2025).
Traditional receivers adapt continuously to SNR variations in filtering, demodulation, and decoding, with no discrete thresholds separating processing regimes (Tse and Viswanath (2005); Goldsmith (2005)). The continuous SNR-driven manifold observed in the activation space reflects this, suggesting the model has internalized the gradual nature of signal degradation. This alignment with classical design principles supports the interpretability and physical plausibility of the learned representations.
4.4 Evaluating Clustering Validity by using Out-of-Distribution Detection Task
We aimed to maximize the number of clustered samples, but a high clustering rate is only meaningful when paired with high clustering quality Campello et al. (2013). As this is exploratory analysis, ground truth labels are unavailable, and internal metrics like DBCV or persistence reflect algorithmic performance, not practical utility. Cross-method comparisons are limited by differing preprocessing and metrics. As noted by Jain (2010), clustering utility is domain-specific and application-dependent.
We applied the clustering structure to an Out-of-Distribution (OOD) detection task, assuming OOD samples would fall outside the clustering. Results in Figure 8 support this: the baseline clustered 15.6% of OOD samples; the proposed method, nearly none (2 samples). On In-Distribution (ID) test data, the proposed method had a lower clustering rate (79.0%) than the baseline (91.8%), suggesting better uncertainty handling and less rigid generalization. In other words, it avoids overconfidently clustering unfamiliar patterns, leading to stronger separation between ID and OOD samples. This serves as task-specific validation of clustering quality. The detection task is also practically motivated: results from over-the-air validation (Luostari et al. (2025)) suggest that general-purpose receivers must be able to detect and adapt to distribution mismatch. Further evaluation across diverse scenarios is needed to confirm broader effectiveness.
Figure 7: UMAP embeddings of DeepRx model activations from layer B1-PRE using the baseline (left) and proposed (right) methodology, colored by channel parameters and clustering partitions with HDBSCAN clustering (EoM, minimum cluster size 5). UMAP is used solely for visualization.
Figure 8: UMAP visualization of DeepRx model activations from layer B1-PRE, clustered using HDBSCAN (EoM, minimum cluster size 5). UMAP is used solely for visualization. Circles denote training samples used for clustering; crosses are test samples assigned via HDBSCAN’s predict. Left: baseline methodology; right: proposed methodology. Top: OOD test data (CDL-A, 1000 ns delay spread); bottom: ID test data (CDL channels within DeepRx training range). Visualization illustrates the extent to which ID and OOD test samples fall within or outside the learned clusters.
5 Related Work
5.1 Concept Discovery in Neural Networks
Neuron-centric methods analyze individual neurons to identify features that excite specific units. Techniques include saliency maps (Simonyan et al. (2014)), which highlight input features relevant to activations; neuron maximization (Erhan et al. (2009)), which generates inputs that strongly activate specific neurons; activation histograms (Yosinski et al. (2015)), which summarize neuron behaviors; and input perturbation testing (Fong and Vedaldi (2017)), which evaluates the effects of input alterations on activations or outputs. Grad-CAM (Selvaraju et al. (2017)) produces class-specific visualizations by localizing important regions but focuses on localized activations. Network Dissection (Bau et al. (2017)) aligns neuron behaviors with predefined semantic categories (e.g., textures, objects) but fails to account for distributed patterns across layers.
Concept-based methods take a higher-level perspective by linking input features to human-interpretable concepts (Yeh et al. (2022)). Testing with Concept Activation Vectors (TCAV) (Kim et al. (2018)) measures how user-defined concepts (e.g., stripes, wheels) influence predictions, while automated approaches (Ghorbani et al. (2019)) reduce the need for manual definition. These methods have been extended to tabular data by Pendyala and Choi (2022), but they rely on predefined or discovered concept sets and may miss layer-wide patterns or concept interactions.
Modern deep networks often learn distributed and poly-semantic features, where neurons respond to multiple unrelated stimuli or encode several sub-features (Olah et al. (2017, 2020); Bau et al. (2017, 2020)). This highlights the limitations of focusing on individual neurons or predefined concepts, as individual units may learn similar or no new concepts, or interact with each other to learn a concept (Kim et al. (2018); Ghorbani et al. (2019); Bäuerle et al. (2022)).
Neural Activation Pattern (NAP) methodology (Bäuerle et al. (2022)) addresses these limitations by identifying groups of inputs with similar activation profiles across a layer, capturing layer-level concepts without retraining the model. It clusters activation patterns to reveal distributed features and visualizes them via the NAP Magnifying Glass, which presents representative images, activation statistics, and metadata filtering for comparison across layers and models, as demonstrated in a visual object recognition task by Bäuerle et al. (2022). NAP complements other interpretability methods and supports use cases such as TCAV-based testing or ante-hoc training, where interpretability is integrated during model development (Gupta et al. (2023)). By leveraging predefined concepts, NAP has the potential to enhance both interpretability and model performance.
5.2 Clustering Methods for Concept Discovery
Clustering is an unsupervised learning task that partitions data such that instances within the same cluster are more similar to each other than to those in different clusters (Jain and Dubes (1988)). Selecting an effective clustering method is essential for concept discovery, as it plays a critical role in uncovering patterns within high-dimensional activation spaces.
Hierarchical clustering creates a tree-like partition hierarchy. Methods can be categorized into agglomerative (bottom-up; e.g., BIRCH Zhang et al. (1996)) and divisive (top-down; e.g., PNN Virmajoki (2004)) approaches. Agglomerative methods iteratively merge clusters based on linkage criteria such as single linkage (shortest distance), complete linkage (farthest distance), average linkage (mean distance), or Ward’s method (minimizing variance increase). Divisive methods recursively split clusters and can be monothetic (splitting by a single attribute) or polythetic (multiple attributes).
Partitional clustering divides data into non-overlapping clusters and methods can be categorized based on the framework proposed by Tuononen (2006). This categorization includes optimization-based methods (e.g., K-Means Steinhaus (1956); MacQueen (1967); Lloyd (1982)) that minimize an objective function such as intra-cluster variance; graph-theoretic methods (e.g., spectral clustering Ng et al. (2001)) that partition similarity graphs; density-based methods (e.g., DBSCAN Ester et al. (1996)) that identify dense regions; grid-based methods, which discretize the feature space; and model-based methods (e.g., Gaussian Mixture Models Dempster et al. (1977), Self-Organizing Maps Kohonen (1990)) that fit statistical or neural network models.
HDBSCAN (Campello et al. (2013)) combines density-based and hierarchical clustering by using a minimum spanning tree for single-linkage agglomeration, effectively handling noise and irregular boundaries—ideal for high-dimensional data like neural activations with unknown cluster counts. Its accelerated variant, HDBSCAN* (McInnes and Healy (2017)), improves efficiency through optimized neighbor searches and pruning, making it well-suited for large-scale clustering tasks. The NAP framework (Bäuerle et al. (2022)) adopts this approach to reveal complex activation structures.
5.3 Evaluation of Cluster Validity
Evaluating cluster validity is essential in clustering-based concept discovery, as it determines whether the identified groupings reflect meaningful structure in the data. While many validation methods exist (Hassan et al. (2024)), choosing the right one is nontrivial: the utility of a clustering often depends highly on the specific task and domain. As Jain (2010) noted, clustering is in the eye of the beholder.
Cluster validation criteria fall into three categories (Jain and Dubes (1988)): external (compare to ground truth labels, e.g., Rand index Rand (1971), Jaccard index Jaccard (1901)); internal (assess intrinsic structure, e.g., Silhouette score Rousseeuw (1987), Davies–Bouldin score Davies and Bouldin (1979)); and relative (compare multiple clustering results, e.g. Halkidi et al. (2001)). Key indicators of cluster quality are (Jain and Dubes (1988)) compactness (minimizing distances within clusters) and separation (maximizing distances between clusters). For density-based clustering algorithms, internal validation criteria should favor high-density clusters with low-density separation. Two widely used internal indices for density-based clustering are CDbw (Halkidi and Vazirgiannis (2008)), which uses boundary points to assess structure, and DBCV (Moulavi et al. (2014)), preferred for its more accurate density estimation and boundary-free formulation.
6 Conclusions
We evaluated the Neural Activation Pattern (NAP) methodology for discovering layer-level concepts by clustering full-layer activation distributions. In both visual object recognition and radio receiver experiments, the baseline methodology with recommended HDBSCAN parameters resulted in a majority of samples remaining unclustered, with rates exceeding 90% in some cases as data volume increased. The use of Excess of Mass (EoM) as a cluster selection criterion effectively reduced this issue, while our proposed improvements—spanning normalization, distribution estimation, and distance metrics—led to more stable clustering performance across configurations.
Although distinct layer-level concepts were not identified, a continuous manifold shaped by Signal-to-Noise Ratio (SNR) consistently emerged in 2D UMAP embeddings. This structure aligns with classical receiver behavior, suggesting physical plausibility and highlighting SNR as a dominant learned factor in the receiver model’s internal representation. Our proposed methodology also improved separation between In-Distribution (ID) and Out-of-Distribution (OOD) samples, indicating stronger generalization and supporting OOD detection as both a practical application—motivated by real-world deployment needs—and an indirect measure of clustering quality. Limitations include the lack of analysis of variability due to neural network initialization and training stability, as well as the limited evaluation of OOD detection, which was tested in only a single scenario.
These findings suggest several directions for future work. The SNR-shaped manifold could support improved OOD detection through SNR estimation in receiver systems and nearest-neighbor reasoning in activation space, following runtime monitoring approaches such as Cheng et al. (2019). Additionally, since single-layer analysis did not reveal concepts for other channel parameters (e.g., delay spread), exploring concept emergence across multiple successive layers could uncover more complex structures—potentially guided by the Information Bottleneck framework by Shwartz-Ziv (2022).
Overall, our results highlight the importance of robust clustering preprocessing and parameter selection for concept discovery, with practical implications for interpretability, model troubleshooting, and runtime monitoring in deep neural networks—particularly in safety-critical, high-stakes settings such as 6G physical-layer wireless communication.
Acknowledgments and Disclosure of Funding
Ville Hautamäki was partially supported by Jane and Aatos Erkko Foundation. The authors would like to thank the creators of DeepRx and NAP framework for their foundational work. The authors would also like to thank Mikko Honkala and Heikki Penttinen for their contributions.
Supplementary Material
[…]
Appendix A Baseline Neural Activation Pattern (NAP) Methodology
The Neural Activation Pattern (NAP) methodology (Bäuerle et al. (2022)) addresses the limitations of focusing on individual neurons or predefined concepts—as described in Section 5—by identifying groups of inputs with similar activation profiles across a neural network layer. As a model-specific method, it relies on inner model details and provides post-hoc layer-level concept-based interpretability. The key steps in the NAP methodology are as follows:
  • Location Disentanglement: For convolutional layers, spatial information in activation values can act as a confounding factor (Bäuerle et al. (2022)). To address this, activations of each convolutional channel are aggregated using statistical measurements, either maximum value (highest activation), minimum and maximum values (range of activations), mean value (average amount of activations), or mean and standard deviation (average amount and spread of activations). The disentangled activations are denoted as $\tilde{A}^{l}=\{\tilde{a}^{l}_{i}\mid i\in[1,\ldots,N]\}$. Aggregating activations removes spatial information and reduces computational cost, making the method scalable for larger models.
  • Clustering Analysis: The aggregated and normalized activation outputs $\hat{A}^{l}$ are clustered, forming a clustering $C^{l}=\{c^{l}_{j}\mid j\in[1,\ldots,M]\}$ with $M$ clusters, using the HDBSCAN* (McInnes and Healy (2017)). The clustering process employs the Euclidean distance metric, selects clusters from the leaves of the cluster tree, and uses the minimum cluster size of five. Each cluster $c^{l}_{j}\in C^{l}$ groups together a set of inputs $x_{i}\in X$ which produce similar activations for layer $l$. Clusters are mutually exclusive but not collectively exhaustive, meaning that no input $x_{i}$ can belong to more than one cluster, but some inputs $x_{i}$ may not belong to any cluster (i.e., they are classified as noise).
  • Post-Clustering Analysis: HDBSCAN* provides information about the persistence of the clusters. Persistence score reflects the stability or strength of a cluster; higher persistence meaning cluster being more well-defined and robust. Clusters are sorted based on their persistence scores, which help identify the most stable and reliable clusters. Clusters are considered to be just "raw mathematical groupings," while NAPs in general are clusters that reflect features learned by the network. Concepts are interpretable NAPs that align with human understanding.
Appendix B Neural Activation Pattern (NAP) Methodology Implementation Details
Although Bäuerle et al. (2022) mention the availability of their NAP methodology as a Python package, we were unable to locate a public implementation. Consequently, we reimplemented both their baseline approach (Appendix A) and our proposed enhancements (Section 2) using various Python libraries from which the most essential can be found from Listing 5.
In our methodology, the density is estimated using a Kernel Density Estimator (KDE) with a Gaussian kernel $K$. To obtain a compact yet informative representation, the estimated probability density function $\hat{f}^{l}_{i,u}$ (Equation 2) is sampled at $R=10$ evenly spaced Z-scores (Equation 5) over the interval $[z_{\text{min}},z_{\text{max}}]=[-2.576,2.576]$, which contains $P=99\%$ of the mass of a standard normal distribution (Equation 4). This one-point-per-decile grid (step $\Delta\approx 0.515$) yields a ten-dimensional feature vector per unit—an order-of-magnitude increase in information over a single-mean summary in baseline methodology, while remaining compact enough for downstream modeling or tabular presentation.
Appendix C Visual Object Recognition Model (ResNet50 with Pretrained Weights) and Data (ImageNet ILSVRC2012)
We employed a ResNet-50 convolutional neural network pre-trained on the ImageNet-1K dataset—namely IMAGENET1K_V2 weights from TorchVision (Marcel and Rodriguez (2010))—to perform visual object classification. To align with the experimental protocol of Bäuerle et al. (2022), we used a randomly selected 1% subset (12,832 images) of the ImageNet ILSVRC2012 training set (Russakovsky et al. (2015)). As part of the setup—and following their methodology—we extracted feature activations from the Conv4 and Conv5 stages of the network—corresponding to the layer3 and layer4 modules in TorchVision’s implementation—which produce intermediate activations of size [1024 × 14 × 14] and [2048 × 7 × 7], respectively, when processing 224 × 224 resolution images.
Appendix D Radio Receiver Model and Data (DeepRx Trained with Simulated Data)
DeepRx (Honkala et al. (2021)) is a fully convolutional deep learning model that jointly performs the tasks of channel estimation, equalization, and demapping, as illustrated in Figure 9. Moreover, the model is designed to be compliant with 5G NR specifications, capable of estimating the log-likelihood ratio (LLR), or the uncertainty, of each received bit so that it can be fed into an LDPC decoder. The model supports different demodulation reference signal (DMRS) or pilot configurations included in the 5G NR specifications (Honkala et al. (2021)).
Figure 9: Illustration of the PUSCH simulator (adapted from Honkala et al. (2021)), where DeepRx is used to replace channel estimation, equalization, and demapping blocks.
D.1 Data Simulation
The data are randomly generated using a 5G-compliant physical uplink shared channel (PUSCH) simulator, powered by NVIDIA Sionna (Hoydis et al. (2023)), and configured with the parameters listed in Table 1. The simulation process is presented in Figure 9 and consists of three main stages:
Table 1: Simulation parameters used for DeepRx training, as well as NAP training and testing. In-Distribution (ID) testing for NAP uses parameters similar to those used during NAP training. Out-of-Distribution (OOD) testing uses a fixed 1000 ns delay spread that exceeds the range seen during DeepRx training but remains within realistic bounds—specifically, corresponding to a very long delay spread as defined in Table 7.7.3-1 of ETSI (2020).
Parameter DeepRx training NAP training NAP testing (OOD)
and testing (ID)
@@ D.1 Data Simulation
Signal-to-Noise Ratio (SNR) -10 dB to 25 dB -10 dB to 25 dB -10 dB to 25 dB
Doppler Shift 0 Hz to 330 Hz 0 Hz to 300 Hz 0 Hz to 300 Hz
Delay Spread 10 ns to 300 ns 10 ns to 300 ns Fixed 1000 ns
Channel Models UMi, UMa CDL-A, E for train CDL-A
All CDL for test

### v2

@@ 3 Methodology
where $\mu_{\ell,c}$ and $\sigma_{\ell,c}$ are computed over all spatial locations and training samples and $K$ is a Gaussian kernel with bandwidth $h_{\ell,c}$. We use Scott’s Rule scott1992multivariate for selecting the bandwidth, providing a standard bias-variance tradeoff:
$\phantom{.}h_{\ell,c}=1/\sqrt[5]{H_{\ell}W_{\ell}}.$ (10)
To obtain fixed-length features, we sample the KDE over a Z-score interval covering $P\%$ of activation mass,
$z_{\min},z_{\max}]=\left[\Phi^{-1}\!\Bigl(\tfrac{1-P/100}{2}\Bigr),\Phi^{-1}\!\Bigl(\tfrac{1+P/100}{2}\Bigr)\right],$ (11)
where $\Phi^{-1}(\cdot)$ is inverse cumulative distribution function of the standard normal distribution bliss1934probit, using $R$ uniformly spaced points,
$\hat{p}_{\ell,i,c}^{(r)}=\hat{p}_{\ell,i,c}\!\left(z_{\min}+r\,\frac{z_{\max}-z_{\min}}{R-1}\right),$ (12)
where $r\in[0,R-1]$. The per-channel KDE vector is
$\mathbf{k}_{\ell,i,c}=[\hat{p}_{\ell,i,c}^{(0)},\dots,\hat{p}_{\ell,i,c}^{(R-1)}],$ (13)
and concatenation across channels yields
$\mathbf{k}_{\ell,i}=[\mathbf{k}_{\ell,i,1},\dots,\mathbf{k}_{\ell,i,C_{\ell}}].$ (14)
(iii) Free-energy features.
The absolute energy for channel $c$ is computed as liu2020energy
$e_{\ell,i,c}=-\log\sum_{q=1}^{H_{\ell}W_{\ell}}\exp\!\bigl(\mathbf{a}_{\ell,i,c}[q]\bigr),$ (15)
and concatenation across channels yields
$\mathbf{e}_{\ell,i}=[e_{\ell,i,1},\dots,e_{\ell,i,C_{\ell}}].$ (16)
SNR-Based Complementary Features. Let $p_{i}\in\mathbb{R}$ denote the estimated Signal-to-Noise-Ratio (SNR) of input $x_{i}$. SNR estimate $p_{i}$ can be used directly as an auxilary feature or through the SNR distance
$d_{\mathrm{SNR}}(i,j)=|p_{i}-p_{j}|.$ (17)
To better separate samples that are close in activation space but correspond to different channel conditions, we incorporate SNR into OOD detectors. In particular, we use conditional k-NN schemes, where neighbor selection is either restricted to samples with similar SNR (hard conditioning) or smoothly weighted by SNR mismatch (soft conditioning), analogous to stratified k-NN katila2002 but with a continuous auxiliary variable. Full definitions are provided in Supplemental E.
OOD Modeling. Let $\mathbf{g}_{\ell,i}$ denote the feature for layer $\ell$ and sample $i$. OOD detectors are fitted on the In-Distribution (ID) training set and produce a scalar OOD score. HDBSCAN and k-NN rely on an explicit distance metric $d(\cdot,\cdot)$, whereas OC-SVM and Mahalanobis induce distances implicitly via a learned decision boundary or covariance structure.
(i) Density-based (HDBSCAN).
We cluster ID features using HDBSCAN* campello2013hdbscan; mcinnes2017accelerate and use the cluster-membership strength
$s_{\ell,i}^{\mathrm{HDBSCAN}}=1-\mathrm{membership}\bigl(\mathbf{g}_{\ell,i}\bigr),$ (18)
since low membership indicates noise and is treated as OOD.
(ii) Neighbourhood-based (k-NN).
We train on ID features and use the mean distance to the $k$-nearest neighbours cover1967,
$s_{\ell,i}^{\mathrm{kNN}}=\frac{1}{k}\sum_{j\in\mathcal{N}_{k}(i)}d(\mathbf{g}_{\ell,i},\mathbf{g}_{\ell,j}),$ (19)
where $\mathcal{N}_{k}(i)$ denotes the $k$-nearest neighbours in training set for sample $i$. We do not apply Deep $k$-Nearest Neighbors papernot2018deepknn, as it relies on class-label agreement, which is undefined for multi-label soft-bit outputs.
(iii) Boundary-based (One-Class SVM).
We train on ID features and use the distance to decision boundary scholkopf2001ocsvm; bounsiar2014ocsvm,
$s_{\ell,i}^{\mathrm{OCSVM}}=-b_{\ell}(\mathbf{g}_{\ell,i}),$ (20)
where $b_{\ell}(\mathbf{g}_{\ell,i})=\sum_{j}\alpha_{\ell,j}K(\mathbf{g}_{\ell,i},\mathbf{g}_{\ell,j})-\rho$ is the learned decision function (positive inside the ID region) with kernel $K$.
(iv) Covariance-based (Gaussian Mahalanobis).
Assuming Gaussian ID features,
$s_{\ell,i}^{\mathrm{Mahalanobis}}=\bigl(\mathbf{g}_{\ell,i}-\mu_{\ell}\bigr)^{\top}\Sigma_{\ell}^{-1}\bigl(\mathbf{g}_{\ell,i}-\mu_{\ell}\bigr),$ (21)
where $(\mu_{\ell},\Sigma_{\ell})$ are the empirical mean and covariance mahalanobis1936generalised. We do not adopt the class-conditional Mahalanobis variants of lee2018mahalanobis; muller2025mahalanobis, as they rely on per-class distributions, which are ill-defined for multi-label receiver outputs.
Fusion Across Detectors and Layers (Optional). Each layer $\ell$ yields a set of scores
$\mathbf{s}_{\ell,i}=\bigl[s_{\ell,i}^{\text{kNN}},s_{\ell,i}^{\text{Mahalanobis}},s_{\ell,i}^{\text{HDBSCAN}},s_{\ell,i}^{\text{OCSVM}},p_{i}\bigr].$ (22)
Within-layer fusion combines (two or more) detector outputs at a single layer, either by fitting an OC-SVM on scores or by converting the scores to conformal $p$-values and applying a Benjamini–Hochberg (BH) benjamini1995bh style combination rule magesh2023conformalpvalues. Within-layer fusion leverages classifier diversity.
Across-layer fusion combines layer-level scores (of one or more detectors) either by fitting OC-SVM or applying BH style combining, yielding a continuous OOD score via Supplemental F. Layer fusion leverages across-layer diversity.
OOD Decision. Let $S_{i}$ denote the final OOD score for sample $x_{i}$. A binary decision is obtained by thresholding,
$x_{i}\text{ is OOD}\quad\Longleftrightarrow\quad S_{i}\geq\tau,$ (23)
where the threshold $\tau$ is chosen to satisfy a desired operating point (e.g., FPR@95%TPR). The continuous score $S_{i}$ can also be used to generate Detection Error Tradeoff (DET) curves martin1997det.
Our framework is post-hoc: the receiver parameters $\theta$ are kept frozen. It is not model-agnostic, as feature extraction assumes convolutional layers with channelwise structure, as in ResNet-like neural receivers such as DeepRx honkala2021deeprx. The block diagram in Fig. 2 depicts the full dataflow.
Table 1: Simulation parameters for DeepRx and Out‑Of‑Distribution (OOD) detector training/testing. In‑Distribution (ID) testing mirrors the OOD‑training parameters. OOD scenarios: (i) fixed 1000 ns delay spread (very long multipath; Table 7.7.3‑1 of 3gpp38901), (ii) fixed 866 Hz Doppler (approximately 100 m/s at 2.6 GHz; high‑speed rail). These OOD settings exceed DeepRx’s training range yet remain realistic. OOD testing uses the Clustered Delay Line (CDL)‑C model to induce a channel‑model mismatch; training uses CDL‑A and CDL‑E.
Parameter DeepRx train Detector train Detector OOD test Detector OOD test
and ID test High DS High Speed
@@
Signal-to-Noise Ratio (SNR) -10–25 dB -10–25 dB -10–25 dB -10–25 dB
Doppler Shift 0–330 Hz 0–300 Hz 0–300 Hz 866 Hz
Delay Spread 10–300 ns 10–300 ns 1000 ns 10–300 ns
Channel Models UMi, UMa CDL-A, E (train) CDL-C CDL-C
CDL-C (test)